OpenStudy (anonymous):

Put the linear equation into standard form: y=mx+b (-1/2)x+(2/3)y=10

OpenStudy (anonymous):

-0.5x+(2/3)y=10 add 0.5x on both sides -0.5x+0.5x+(2/3)y=0.5x+10 (2/3)y=0.5x+10 divide both sides by (2/3) y = (3/4)x+15